A cylinder has an applied pressure of 600 torr and a volume of 7.0 L. What is the volume if the applied pressure increases to 5 atm?
Answer #1

Given values are-

P1 = 600 torr = (600÷ 760) atm = 0.7895 atm

P2 = 5 atm, V1 = 7.0 L , V2 = ?

=> From Boyles law, (at constant temperature)

P1V1 = P2V2

0.7895 × 7.0 = 5 × V2

=> 5.5265 = 5 × V2

=> V2 = 5.5265 ÷ 5 = 1.1053 L

The volume becomes 1.1053 L when pressure increases to 5 atm.